Appendix I

Miscellaneous Exercises

[56/57%]
Show
LaTeX
Problem I-1

For any two events AA, BB show that the following relationships are equivalent: A⊂BA \subset B, A′⊃B′A' \supset B', A∪B=BA \cup B = B, and A∩B′=∅A \cap B' = \emptyset.

?
Problem I-2

Show that (A△B)′=(AB)∪(A′B′)(A \triangle B)' = (AB) \cup (A'B') and P(A△B)=P(A)+P(B)−2P(AB)\mathbb {P}\left(A \triangle B\right) = \mathbb {P}\left(A\right) + \mathbb {P}\left(B\right) - 2\mathbb {P}\left(AB\right), where △\triangle is the symmetric difference, defined in Exercise 34.

?
Problem I-3

Let nn be the number of elements in the sample space Ω\Omega and N(E)N(E) the number of elements in the event EE. Show that the data: N=1,000N = 1{,}000, N(A)=525N(A) = 525, N(B)=312N(B) = 312, N(C)=470N(C) = 470, N(AB)=42N(AB) = 42, N(AC)=147N(AC) = 147, N(BC)=86N(BC) = 86, and N(ABC)=25N(ABC) = 25 are inconsistent.

?
Problem I-4

(Lewis Caroll, A Tangled Tale, 1881). In a fierce battle, at least 70 of the soldiers lost an eye, at least 74 lost an ear, at least 80 lost an arm, and at least 85 lost a leg. What is the smallest possible percentage of soldiers that each lost an eye, an ear, an arm, and a leg?

?
Problem I-5

In a market research survey, from a sample of 1,000 smokers, 811 said that they prefer brand $A$'', 752 prefer BB'', and 418 prefer $C$''. Furthermore, 570 said that they like both AA'' and $B$'', 356 that they like both AA'' and $C$'', 348 that they like both BB'' and ``CC'', and 297 that they like all three brands. How did the researcher discover that the results were inconsistent?

?
Problem I-6

We throw two dice once. Let AA: doubles, BB: at least one die shows 5 or 6, CC: the sum of the outcomes of the two dice is less than 6.

?
(a)

Find P(A)\mathbb {P}\left(A\right), P(B)\mathbb {P}\left(B\right), P(C)\mathbb {P}\left(C\right), P(AB)\mathbb {P}\left(AB\right), and P(AC)\mathbb {P}\left(AC\right).

(b)

Using only the results of (a) calculate: P(A∪B)\mathbb {P}\left(A \cup B\right), P(B∪C)\mathbb {P}\left(B \cup C\right), and P(A∪C)\mathbb {P}\left(A \cup C\right).

Problem I-7

Three athletes are participating in the Marathon. Their chances of finishing under 2 hours are 1/31/3, 1/51/5, and 1/121/12, respectively. Their times for the duration of the course are independent. What is the probability that at least one finishes in less than 2 hours?

?
Problem I-8

A horse has probability pp of jumping over a specific obstacle. Given that in 5 trials it succeeded 3 times, what is the conditional probability that it had succeeded in the first trial?

?
Problem I-9

In a train car 3 seats are facing the front and 3 are facing the rear. Two women and three men enter the car and sit at random. What is the probability:

?
(a)

that the two women are facing each other?

(b)

that two men are facing each other?

Problem I-10

On the basis of the following data, examine the relationship between the father having brown eyes (event AA) and the son having brown eyes (event BB). The percentages observed were: ABAB: 5, AB′AB': 7.9, A′BA'B: 8.9, and A′B′A'B': 78.2.

?
Problem I-11

We throw three dice. What is the probability that at least one 6 appears given that the three outcomes were completely different. Generalize the result for the case of a regular polyhedral die.

?
Problem I-12

Ninety percent of a population have a television set (event AA), while 80 have a car (event BB). What is the minimum percentage of the population with a car that also have a television?

?
Problem I-13

Let P(A)=α\mathbb {P}\left(A\right) = \alpha and P(B)=1−ε\mathbb {P}\left(B\right) = 1 - \varepsilon where ε\varepsilon is a small number. Give the upper and lower bound of P(A∣B)\mathbb {P}\left( A \mid B \right).

?
Problem I-14

Let X1,X2X_{1}, X_{2} be independent random Bernoulli variables with the same parameter pp and let Y=0Y = 0 if X1+X2=X_{1} + X_{2} = even and Y=1Y = 1 if X1+X2=X_{1} + X_{2} = odd.

?
(a)

For what values of pp are X1X_{1} and YY independent?

(b)

Using (a), define three events AA, BB, and CC such that they are pairwise independent but not completely independent.

Problem I-15

We throw three dice twice. What is the probability of having the same result the second time when the dice are, (a) not distinguishable? (b) distinguishable?

?
Problem I-16

At a party, AA is introduced to 6 women and 4 men. What is the probability that he identifies all the couples correctly if he knows that there are, (a) 4 couples, (b) 3 couples?

?
Problem I-17

A young person driving a motorcycle on a highway with 4 lanes changes lane every 15 seconds. What is the probability pp of finding himself in the lane in which he started 1 minute later?

?
Problem I-18

In city AA it rains half of the days of the year and the weather forcast is correct with probability 2/32/3. Mr. Sugar worrying about getting wet always takes his umbrella when the forcast calls for rain and with probability 1/31/3 otherwise. Find the probability that:

?
(a)

he is caught in the rain without his umbrella;

(b)

he has his umbrella with him when there is no rain.

Problem I-19

The telegram signals (⋅)(\cdot ) and (−)(-) are sent with proportion 3 ⁣: ⁣43\! :\! 4. Due to faults in transmission a dot (⋅)(\cdot ) becomes (−)(-) with probability 1/31/3. If a signal has arrived as (⋅)(\cdot ) what is the probability that it was transmitted correctly?

?
Problem I-20

A die is thrown 10 times and let the maximum outcome be MM and the minimum mm. Find the P(m=2,M=5)\mathbb {P}\left(m = 2, M = 5\right).

?
Problem I-21

An urn contains bb black balls and ww white balls. nn balls are drawn from the urn at random, their color unnoticed, then mm additional balls are drawn. Show that the probability that there are kk black balls among the mm balls is the same as if the mm balls were chosen from the beginning (i.e., before choosing the nn balls).

In other words, in sampling from a finite population the final (hypergeometric) probabilities are not affected by the content of an initial sample when we do not know that content. Obviously, if we know the composition of the first sample (i.e., the colors of the nn balls), then, in general, the probabilities are affected.

?
Problem I-22

After an earthquake the books of a large library fall on the floor and the servant, who does not know how to read, reshelves them in any order.

?
(a)

What is the probability that exactly kk books are placed back in their initial position?

(b)

How many books are expected to be reshelved in their initial position?

Problem I-23

John Smith in 1693 posed the following problem: Is the probability for each of the three players to win the same if: the first must obtain at least one ace throwing 6 dice, the second at least two aces throwing 12 dice, and the third at least three aces throwing 18 dice? Newton and Tollet found that the first player has a better chance than the second one, and the second player has a better chance than the third one.

?
Problem I-24

A homogeneous and regular die with vv faces is thrown nn times. Let p(n,k)p(n, k) be the probability that a given face appears less than kk times. Show that:

?
(a)

p(vn,n)p(vn, n) is a decreasing function of vv for given nn;

(b)

p(vn,n)<1/2p(vn, n) < 1/2;

(c)

p(2n,n)→1/2p(2n, n) \rightarrow 1/2 as n→∞n \rightarrow \infty (De Moivre central limit theorem).

Problem I-25

The Massachusetts State Lottery issues 1 million tickets. The first winning ticket wins $50,000, the next 9 winning tickets win $2,500 each, the next 90 tickets win $250 each, and 900 tickets win $25 each. What is the expected gain of a person buying, (a) one ticket? (b) ten tickets?

?
(a)
(b)
Problem I-26

(Continuation of Problem I-25.) What is the answer to (a) and (b) if:

  1. the state issues 2 million tickets?

  2. the winning tickets win twice as much?

?
Problem I-27

In a survey on the housing problem, 2 apartments are chosen from each of 4 buildings (having both owner-occupied as well as rented apartments). If we know that:

Building No.1234
Owner-occupied5482
Rented810910

What is the expected number of owner-occupied apartments in the above sample?

?
Problem I-28

Out of 365 people:

?
(a)

What is the probability that at least two have their birthday on 2 April?

(b)

What is the expected number of people having their birthday on 2 April?

(c)

What is the expected number of days in a year that are birthdays:

  1. for at least one person?

  2. for at least two persons?

How can you approximate the answers to the above questions?

Problem I-29

The duration TT of a telephone call satisfies the following equation:

P(T>t)=p⋅exp⁡[−λt]+(1−p)⋅exp⁡[−μt]for t>0, \mathbb {P}\left(T > t\right) = p \cdot \exp [-\lambda t] + (1 - p) \cdot \exp [-\mu t] \qquad \text{for } t > 0,

where 0≤p≤10 \leq p \leq 1, λ>0\lambda > 0, μ>0\mu > 0. Find E[T]\mathbb {E}\left[T\right] and Var⁡[T]\operatorname {Var}\left[T\right].

?
Problem I-30

Two gamblers AA and BB play a series of independent games in each of which AA has probability pp of winning and BB has probability q=1−pq = 1 - p of winning. Suppose that at a given moment the series of games is interrupted and AA needs mm additional points and BB needs nn additional points to be the winner. The bet should be split between AA and BB in the ratio P(A):P(B)\mathbb {P}\left(A\right):\mathbb {P}\left(B\right), where P(A)=1−P(B)\mathbb {P}\left(A\right) = 1 - \mathbb {P}\left(B\right) is the probability that AA eventually wins. Show that

PA=∑k=mm+n−1(m+n−1k)pk−qm+n−1−k=∑m=0n−1(m+k−1k)pmqk. P_{A} = \sum _{k=m}^{m+n-1} \binom {m+n-1}{k} p^{k} - q^{m+n-1-k} = \sum _{m=0}^{n-1} \binom {m+k-1}{k} p^{m} q^{k}.

Note: The above solution was first obtained by Montmort (1675--1719). Explain the idea behind each term in the expressions for PAP_{A}. Try to show the equality between those two expressions for PAP_{A} using the properties of binomial coefficients (see Chapter 1).

?
Problem I-31

An employee leaves between 7:30 a.m. and 8:00 a.m. to go to his office, and he needs 20--30 minutes to get to the railway station. The departure time and the duration of travel are distributed independently and uniformly in the corresponding intervals. He can take either of the following two trains: the first leaves at 8:05 a.m. and takes 35 minutes; and the second leaves at 8:25 a.m. and takes 30 minutes.

?
(a)

What time, on average, does he arrive at his office?

(b)

What is the probability that he misses both trains?

(c)

How many times is he expected to miss both trains in 240 working days?

Problem I-32

Show that if XX and YY are independent, then (cf. Exercise 328)

Var⁡[XY]=Var⁡[X]⋅Var⁡[Y]+Var⁡[X]⋅[E[Y]]2+Var⁡[Y]⋅[E[X]]2. \operatorname {Var}\left[XY\right] = \operatorname {Var}\left[X\right] \cdot \operatorname {Var}\left[Y\right] + \operatorname {Var}\left[X\right] \cdot [\mathbb {E}\left[Y\right]]^{2} + \operatorname {Var}\left[Y\right] \cdot [\mathbb {E}\left[X\right]]^{2}.

Conclude that if X1,…,XnX_{1}, \ldots , X_{n} are completely independent with means zero, then

Var⁡[∏i=1nXi]=∏i=1nVar⁡[Xi]. \operatorname {Var}\left[\prod _{i=1}^{n} X_{i}\right] = \prod _{i=1}^{n} \operatorname {Var}\left[X_{i}\right].
?
Problem I-33

We toss a fair coin until two consecutive heads or tails appear. What is the probability that an even number of tosses will be required?

?
Problem I-34

ss distinguishable balls are placed randomly into nn different cells. Find the probability that exactly rr cells will be occupied.

?
Problem I-35

A large board is divided into equal squares of side aa. A coin of diameter 2r<a2r < a is thrown randomly onto the board. Find the probability that the coin overlaps:

  1. with exactly one square;

  2. with exactly two squares.

?
Problem I-36

A square of side aa is inscribed into a circle. Find the probability that:

  1. any randomly selected point of the circle will be an interior point of the square.

  2. of eight randomly selected points in the circle, three fall inside the square, two in one of the four remaining parts of the circle, and one in each of the other three parts of the circle.

?
Problem I-37

The distance between two successive cars on a highway follows the exponential distribution with mean 20m20m. Find the probability that there are 90 to 110 cars on a 2-km stretch of the highway.

?
Problem I-38

The number of cars crossing a toll bridge follows the Poisson distribution with an average of 30 cars per minute. Find the probability that in 200 seconds no more than 100 cars will pass through.

?
Problem I-39

(Continuation of Problem I-38.) Establish the relation

λn(n−1)!∫t∞e−λxxn−1 dx=∑k=0n−1e−λt(λt)kk!,(*) \frac{\lambda ^{n}}{(n-1)!} \int _{t}^{\infty } e^{-\lambda x} x^{n-1} \, dx = \sum _{k=0}^{n-1} e^{-\lambda t} \frac{(\lambda t)^{k}}{k!}, \tag {*}

between the Erlang distribution function and the Poisson distribution function.

?
Problem I-40

(Continuation of Problem I-39.) Show that

1(n−1)!∫λ∞e−xxn−1 dx=∑k=0n−1e−λλkk!. \frac{1}{(n-1)!} \int _{\lambda }^{\infty } e^{-x} x^{n-1} \, dx = \sum _{k=0}^{n-1} e^{-\lambda } \frac{\lambda ^{k}}{k!}.
?
Problem I-41

Define Φ(a;b)=Φ(b)−Φ(a)\Phi (a; b) = \Phi (b) - \Phi (a) for 0<a<b0 < a < b with Φ\Phi the distribution function of N(0,1)\mathscr {N}\left(0, 1\right). Show that Φ(0;1)>Φ(1;2)\Phi (0; 1) > \Phi (1; 2). Generalize this result for any equal intervals.

?
Problem I-42

Using the fact that if XX follows the normal distribution N(0,1)\mathscr {N}\left(0, 1\right), then

P(X2≤x)=(2/π)1/2∫0xe−u2/2 du, \mathbb {P}\left(X^{2} \leq x\right) = (2/\pi )^{1/2} \int _{0}^{\sqrt{x}} e^{-u^{2}/2} \, du,

conclude that

Γ(1/2)=∫0∞x(1/2)−1e−x dx=π. \Gamma (1/2) = \int _{0}^{\infty } x^{(1/2)-1} e^{-x} \, dx = \sqrt{\pi }.
?
Problem I-43

The probability of hitting a target with one shot is p1p_{1} and the probability of destroying the target with kk (k≥1k \geq 1) shots is 1−p2k1 - p_{2}^{k}. Find the probability of destroying the target after nn shots.

?
Problem I-44

Let XX follow a Poisson distribution with parameter λ\lambda. We select NN random points from the interval [0,1][0, 1]. Let XiX_{i} be the number of points in the interval [(i−1)/n,i/n][(i-1)/n, i/n] (i=1,…,ni = 1, \ldots , n). Show that the random variables XiX_{i} are independent and identically distributed.

?
Problem I-45

If XX and YY are independent with P(X=0)=P(X=1)=1/2\mathbb {P}\left(X = 0\right) = \mathbb {P}\left(X = 1\right) = 1/2 and P(Y≤y)=y\mathbb {P}\left(Y \leq y\right) = y (0≤y≤10 \leq y \leq 1), find the distribution of the following random variables WW, ZZ where:

  1. W=X+YW = X + Y;

  2. Z=XYZ = XY.

?
Problem I-46

Let XX be a uniform random variable on the interval (−a,a)(-a, a) and YY a random variable independent of XX, with distribution function F(y)F(y). Show that the distribution function of G=X+YG = X + Y is

G(z)=12a∫−aaF(z−x) dx. G(z) = \frac{1}{2a} \int _{-a}^{a} F(z - x) \, dx.
?
Problem I-47

Let XX be uniformly distributed on the interval (0,1)(0, 1). Find the probability density function of:

  1. Y=aX+bY = aX + b; a,ba, b constants (a>0a > 0);

  2. Z=1/XZ = 1/X;

  3. W=g(X)W = g(X), where gg is a continuous and monotone function of xx in (0,1)(0, 1).

?
Problem I-48

Each of three players AA, BB, and CC has probability 1/31/3 of hitting a target better than the other two. At each round of the game each shoots at the target once. The winner is the first player to win six rounds. They bet 10 ducats each. At a certain point, however, when AA has 4 victories, BB has 3 victories, and CC has 2 victories, they have to terminate the game. How should they split their bet? (cf. Problem 30).

?
Remark.
?

In addition to Problems I-23 (John Smith, 1963) and I-30 (points of Pascal), it is worth giving here some other problems which are of some historical interest, and show the difficulty in solving probability problems encountered by even well-known figures in the history of mathematics.

Problem I-49

AA and BB play a game in which each of them throws two ordinary dice. AA wins if he scores 6 and BB wins if he scores 7. First, AA throws the dice once, then BB throws the dice twice, and they continue alternating after each throw until one of them wins. What is the probability of winning for each of them?

Note: Huyghens and Spinoza gave the same answer:

P(A)/P(B)=10355/1227≈0.84. \mathbb {P}\left(A\right)/\mathbb {P}\left(B\right) = 10355/1227 \approx 0.84.

(cf. Exercise 67).

?
Remark.
?

He is the author of the first book on probability, De Ratiociniis in Ludo Aleae, 1654.

Problem I-50

A deck of forty cards consists of 10 red, 10 black, 10 blue, and 10 green cards. Four cards are drawn at random and without replacement. What is the probability that all the colors will show up?

(Huyghens gave the answer 1,000/9,129=0.1091{,}000/9{,}129 = 0.109; is it correct?)

?
Problem I-51

An urn contains 8 black balls and 4 white balls. Each of three players AA, BB, and CC (and in that order) draws a ball without replacement and the first one to draw a white ball wins. What is the probability of winning for each of them?

?
Problem I-52

Three players AA, BB, and CC play the following game. They play three rounds for each of which, each of them has probability 1/31/3 of winning. AA is the winner of the game if he wins a round before BB or CC wins two rounds; BB is the winner if he wins two rounds before player AA wins one round, or CC wins two rounds; CC wins like BB. What is the probability of winning for each of them?

(Montmort gave P(A)=15/27\mathbb {P}\left(A\right) = 15/27, P(B)=P(C)=5/27\mathbb {P}\left(B\right) = \mathbb {P}\left(C\right) = 5/27).

?
Problem I-53

A container has 2N2N cards, numbered in such a way that each of the numbers 1,2,…,N1, 2, \ldots , N appears twice. 2N−n2N - n cards are drawn at random. What is the expected number of pairs of cards remaining in the container?

(Bernoulli found n(n−1)/(4N−2)n(n-1)/(4N-2)).

?
Problem I-54

NN lottery tickets are numbered 1 to NN. nn tickets are selected at random. What is the probability that rr consecutive numbers are selected (r=1,2,…,nr = 1, 2, \ldots , n)?

?
Remark.
?

Example: For N=10N = 10, n=4n = 4, r=4r = 4, the probability is 7/(104)=1/307 \big/ \binom {10}{4} = 1/30.

Problem I-55

The probability of event AA is pp. What is the probability that:

?
(a)

in nn trials event AA occurs kk times consecutively?

(b)

in nn trials event AA occurs kk times consecutively and then AcA^{c} also occurs kk times consecutively?

(c)

in n+mn + m trials event AA occurs nn times given that in r+kr + k trials it occurred rr times?

Condorcet gave the following answer for (c):

(r+k+1)! (r+n)! (k+m)!/r! k(r+k+m+n+1)!. (r+k+1)! \, (r+n)! \, (k+m)!/r! \, k(r+k+m+n+1)!.

Laplace and G. Polya disagreed with this solution. What do you say?

Problem I-56

In a lottery with NN tickets, nn numbers are chosen as winners. What is the probability pp that in kk lottery games all the NN numbers will appear?

(Laplace for N=90N = 90, n=5n = 5, and k=85k = 85 gave p=1/2p = 1/2).

?